# Sensing Gravity with Polarized Electromagnetic Radiation

### v1

@@ VII.2 Measuring Angular Momentum of a Gravitational Source using Polarization Wiggling
We will now apply the results for arbitrary scalar and vector perturbations to analyze a more specific case. We will show that the polarization wiggling of electromagentic radiation emitted by a radiation emitter in orbit around a gravitational source with unknown angular momentum can be used to determine the angular momentum of the gravitational source.
@@
where $t$ is time, $(r,\theta,\phi)$ are spherical coordinates, $m$ is the mass of the gravitational source and $j$ is the scalar angular momentum, assuming the rotation axis is the z-axis $\hat{z}$. In Section VI, we found that scalar perturbations do not contribute to polarization wiggling. The only component of the above metric that contributes is the $\mathrm{dt}\mathrm{d\phi}$ term. The vector perturbation is therefore
@@
B×$\text{\boldmath{\hbox{\mathcal{B}}}}\equiv\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}=-\frac{12j}{r^{2}}\mu\sqrt{1-\mu^{2}}\text{\boldmath{\hbox{n}}}\text{\boldmath{\hbox{,}}}$ (50)
@@ VIII Tensor-induced Polarization Wiggling
The tensor perturbation $h_{\mu\nu}^{(T)}$ of eq. (39) is a gauge invariant quantity that is transverse and traceless, conventionally denoted $h_{\mu\nu}^{\mathrm{TT}}$. To give the results of the previous sections a proper demonstration, we will now explore the the polarization wiggle rate $\tilde{\omega}$ arising from a single gravitational tensor mode in linear gravity, assuming a spacetime with a conformally flat background. This will cover spacetimes with both a flat Minkowski background as well as expanding cosmologies.
@@ VIII.1 Wiggle Rate from a Tensor Mode with a Conformally Flat Background
where $h_{s}^{0}$ is a constant of integration. Let us assume that $h_{s}(\eta)$ is varying slowly, with $|h_{s}^{\prime}|\ll|kh_{s}|$. From eq. (57) then follows that
@@
We have now established the time dependency of the polarization amplitude $c_{s}(k,\eta)$ for a general tensor mode. Next, let us evaluate the polarization wiggle rate $\omega$ in the conformal frame.
@@
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\left(e^{-ik\eta}-e^{-ik\mu\eta}\left(\frac{e^{-ik(1-\mu)\eta_{*}}}{{a(\eta_{*})}^{2}}+ik\mu\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\frac{e^{-ik\left(1-\mu\right)\eta^{\prime}}}{{a(\eta^{\prime})}^{2}}\right)\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (64)
Assuming that both the detector and emitter follow rest geodesics, we may assume that the comoving distance between them is constant to lowest order:
@@
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ike^{-ik\eta}\left(1-\frac{e^{i\left(1-\mu\right)k\chi_{*}}}{{a(\eta-\chi_{*})}^{2}}-ik\mu\ \ \operatorname*{\int}\limits_{0}^{\chi_{*}}d\chi\frac{e^{ik\left(1-\mu\right)\chi}}{{a(\eta-\chi)}^{2}}\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (65)
@@
It is apparent from eq. (64) that, if $a(\eta_{*})\ll 1$, the second term on the right-hand side will be the dominant one of the two first terms. This is the case if the polarized radiation and the tensor mode were emitted at early times in an expanding cosmology. In the next two sections, we will evaluate eq. (64) for two different backgrounds: A flat Minkowski background and an expanding, conformally flat background.
@@ VIII.2 Gravitational Polarization Sensitivity $\sigma^{s}$
Next, let us evaluate the polarization sensitivity $\sigma^{s}(\hat{k},\hat{p})$ of eq. (62). It is defined as a purely geometric quantity in terms of two spatial directions $\hat{k}$ and $\hat{p}$ and the tensor polarization basis $\sigma_{\mu\nu}^{s}$. First, the tensor basis $\sigma_{\mu\nu}^{s}(\hat{k})$ for a given direction of propagation ${\hat{k}}^{\mu}$ must be chosen. With the choice of an orthonormal polarization basis $e_{A}^{\mu}(\hat{k}),A=1,2$ perpendicular to the propagation direction $\hat{k}$, the basis tensor $\sigma_{\mu\nu}^{s}$ can be defined by
@@
VIII.3 Wiggle Rate from a Tensor Mode on a Flat Background
Let us assume a flat background with $a(\eta)=\mathrm{const}=1$ and evaluate the wiggle rate of eq. (64). The result is
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\frac{\left(1-e^{i\left(1-\mu\right)k\chi_{*}}\right)}{1-\mu}e^{-ik\eta}\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$
To lowest order, proper time of the observer equals coordinate time; $\tau=\eta+\mathcal{O}(h)$. Thus, integrating with respect to coordnate time $\eta$ gives the polarization angle relative to the mean angle:
$\Omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{\left(\sigma^{+}(\hat{k},\hat{p})h_{+}^{0}(k)+\sigma^{\times}(\hat{k},\hat{p})h_{\times}^{0}(k)\right)\frac{\left(e^{ik(1-\mu)\chi_{*}}-1\right)}{\left(1-\mu\right)}e^{-ik\eta}\right\}.$ (70)
In order to calculate $\Omega^{(T)}$, we need to separate the amplitude and phase of each complex factor on the right-hand side of eq. (70). First, the factor that depends on the distance $\chi_{*}$ to the emitter can be written in polar form as
@@ VIII.3 Wiggle Rate from a Tensor Mode on a Flat Background
$\psi(\mathcal{K},\hat{p})\equiv\arctan(\frac{\sigma_{+}H_{+}\sin\alpha_{+}+\sigma_{\times}H_{\times}\ \ \sin\alpha_{\times}}{\sigma_{+}H_{+}\cos\alpha_{+}+\sigma_{\times}H_{\times}\ \ \cos\alpha_{\times}}).$
@@
$\Omega^{\left(T\right)}(\eta)=W(\mathcal{K},\hat{p},\chi_{*})\sin(k\eta+\gamma(\mathcal{K},\hat{p},\chi_{*})),$ (72)
@@
$W(\mathcal{K},\hat{p},\chi_{*})\equiv 2T(\mathcal{K},\hat{p})\frac{\sin(\frac{\left(1-\mu\right)}{2}k\chi_{*})}{1-\mu}$ (73)
@@
$\gamma(\mathcal{K},\hat{p},\chi_{*})\equiv-\left(\frac{\left(1-\mu\right)}{2}k\chi_{*}+\psi(\mathcal{K},\hat{p})\right).$
The polarization wiggling amplitude $W(\mathcal{K},\hat{p},\chi_{*})$ defines the polarization wiggling response of polarized radiation emitted at distance $\chi_{*}$ from the detector and in direction $\hat{p}$ to a plane gravitational wave with state vector $\mathcal{K}$. The amplitude $W$ and phase $\gamma$ vary with the state $\mathcal{K}$ of the gravitational wave as well as the position of the emitter relative to the detector. We see from eq. (72) that the polarization wiggling frequency is equal to the gravitational wave frequency.
Let us take a look at the conditions under which the amplitude $W$ vanishes. First of all, for inclinations $\mu=\pm 1$, the direction of the emitted radiation is aligned with the direction of the gravitational tensor mode; $\hat{p}=\pm\hat{k}$. From eqs. (68) and (69) follows that $\sigma^{+}(\hat{k},\pm\hat{k})=\sigma^{\times}(\hat{k},\pm\hat{k})=0$, which implies that
$W(\mathcal{K},\pm\hat{k},\chi_{*})=0.$
@@
Next, there are certain emitter distances that give vanishing $W$ for $|\mu|<1$:
$\chi_{*}=\frac{2\pi N}{\left(1-\mu\right)k},N=1,2,...$
With the exception of these cases as well as other border cases where $T(\mathcal{K},\hat{p})$ vanishes for certain combinations of gravitational wave state $\mathcal{K}$ and polarized radiation direction $\hat{p}$, $W(\mathcal{K},\hat{p},\chi_{*})\neq 0$. Therefore, polarized electromagenetic radiation is in general subject to polarization wiggling in the presence of gravitational radiation.
VIII.4 Wiggle Rate from a Tensor Mode in an Expanding Universe
Let us assume a conformally flat background with ${a(\eta)}^{\prime}\neq 0$ and evaluate the wiggle rate of eq. (64). In order to get analytical results and gain an understanding of the results, we will make the simplifying assumption that the observed electromagnetic radiation is emitted in a matter-dominated spacetime. Although not a realistic model of our universe at later times, we expect that this approximation is a reasonable approximation for tensor modes of smaller scale emitted early enough to be sufficiently damped at the end of the matter-dominate era to not yield significant contributions to the wiggling rate from later times. In addition, it is a useful approximation to get a qualitative understanding of how gravitational tensor modes induce polarization wiggling in an expanding universe.
It is known that a matter-dominated universe has a scale factor
$a(\eta)={\left(\frac{\eta}{\eta_{0}}\right)}^{2},$
where the present conformal time is
$\eta_{0}=\frac{2}{H_{0}\sqrt{\Omega_{m}}},$
and the conformal time of emission is
$\eta_{*}=\eta_{0}\sqrt{a(\eta_{*})}.$
$H_{0}$ is the Hubble constant and $\Omega_{m}$ is the present matter fraction.
This enables us to evaluate the third term on the right-hand side of eq. (64). Reparametrizing the integral in terms of the variable $x\equiv k\eta$, we get
$ik\mu\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\frac{e^{-ik\left(1-\mu\right)\eta^{\prime}}}{{a(\eta^{\prime})}^{2}}=ik\mu\eta^{2}J,$
@@ VIII.4 Wiggle Rate from a Tensor Mode in an Expanding Universe
$J\equiv\operatorname*{\int}\limits_{x_{*}}^{x}dt\frac{e^{-i\left(1-\mu\right)t}}{t^{2}}$
and $x_{*}\equiv k\eta_{*}$. The integral $J$ can be expressed in terms of the exponential integral $E_{2}$ as follows [32]:
$J=\frac{E_{2}(i\left(1-\mu\right)x_{*})}{x_{*}}-\frac{E_{2}(i\left(1-\mu\right)x_{0})}{x_{0}}.$
@@
$\tilde{\omega}(\eta)=\mathrm{Re}\left\{ik\left(1+i\mu\right)\left(1-\frac{{e}^{ik(1-\mu)\chi_{*}}}{{a(\eta_{*})}^{2}}\right)e^{-ik\eta}\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}$ (74)
to leading order in $1/k{\eta}_{*}$. This can readily be integrated to obtain the polarization angle $\Omega^{(T)}$ as a function of conformal time:
$\Omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{e^{-ik\eta}(1+i\mu)\left(\frac{{e}^{i\left(1-\mu\right)k\chi_{*}}}{{a(\eta_{*})}^{2}}-1\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (75)
It is evident from eq. (74) that for radiation emitted early, while $a(\eta_{*})\ll 1$, the ${a(\eta_{*})}^{-2}$ terms on the right-hand side of eq. (74) will be dominant, and the wiggle rate depends primarily on the tensor mode amplitude at emission. This confirms our earlier claim that, in an expanding universe, the dominant contributions to the present wiggle rate of polarized radiation emitted at early times come from gravitational fields present at the time of emission. Taking the real part of the complex expression of eq. (75) yields
$\Omega^{\left(T\right)}(\eta)=W(\mathcal{K},\hat{p},\chi_{*})\cos(k\eta+\gamma(\mathcal{K},\hat{p},\chi_{*}))$ (76)
where $T(\mathcal{K},\hat{p})$ is given by eq. (71) above and
$W(\mathcal{K},\hat{p},\chi_{*})\equiv\frac{T(\mathcal{K},\hat{p})\sqrt{1+\mu^{2}}}{{a(\eta_{*})}^{2}}$ (77)
$\gamma(\mathcal{K},\hat{p},\chi_{*})\equiv-\left(\xi+\left(1-\mu\right)k\chi_{*}+\psi(\mathcal{K},\hat{p})\right).$ (78)
We see from eq. (76) that the polarization wigging frequency is equal to the gravitational wave frequency.

### v2

@@ VII.2 Measuring Angular Momentum of a Gravitational Source using Polarization Wiggling
We will now apply the results for arbitrary scalar and vector perturbations to analyze a more specific case. We will show that the polarization wiggling of electromagnetic radiation emitted by a radiation emitter in orbit around a gravitational source with unknown angular momentum can be used to determine the angular momentum of the gravitational source.
@@
where $t$ is time, $(r,\theta,\phi)$ are spherical coordinates, $m$ is the mass of the gravitational source and $j$ is the scalar angular momentum, assuming the rotation axis is the z-axis $\hat{z}$. In Section VI, we found that scalar perturbations do not contribute to polarization wiggling. The only component of the above metric that contributes is the $\mathrm{dt}\mathrm{d\phi}$ term. The vector perturbation is
@@
B×$\text{\boldmath{\hbox{\mathcal{B}}}}\equiv\text{\boldmath{\hbox{\nabla}}}\times\text{\boldmath{\hbox{w}}}=-\frac{12j}{r^{2}}\mu\sqrt{1-\mu^{2}}\text{\boldmath{\hbox{n}}},$ (52)
@@ VIII Tensor-induced Polarization Wiggling
The tensor perturbation $h_{\mu\nu}^{(T)}$ of eq. (39) is a gauge invariant quantity that is transverse and traceless, conventionally denoted $h_{\mu\nu}^{\mathrm{TT}}$. To give the results of the previous sections a proper demonstration, we will now explore the polarization wiggle rate $\tilde{\omega}$ arising from a single gravitational tensor mode in linear gravity, assuming a spacetime with a conformally flat background. This will cover spacetimes with both a flat Minkowski background as well as expanding cosmologies.
@@ VIII.1 Wiggle Rate from a Tensor Mode with a Conformally Flat Background
where $h_{s}^{0}$ is a constant of integration. Let us assume that $h_{s}(\eta)$ is varying slowly with time, with $|h_{s}^{\prime}|\ll|kh_{s}|$. From eq. (59) then follows that
@@
We have now established the time dependency of the polarization amplitude $c_{s}(k,\eta)$ for a tensor mode. Next, let us evaluate the polarization wiggle rate $\omega$ in the conformal frame.
@@
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\left(\frac{e^{-ik\eta}}{{a(\eta)}^{2}}-e^{-ik\mu\eta}\left(\frac{e^{-ik(1-\mu)\eta_{*}}}{{a(\eta_{*})}^{2}}+ik\mu\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\frac{e^{-ik\left(1-\mu\right)\eta^{\prime}}}{{a(\eta^{\prime})}^{2}}\right)\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (66)
Assuming that both the detector and emitter follow rest frame geodesics, we may assume that the comoving distance between them is constant to lowest order:
@@
$\omega^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ike^{-ik\eta}\left(\frac{1}{{a(\eta)}^{2}}-\frac{e^{i\left(1-\mu\right)k\chi_{*}}}{{a(\eta_{*})}^{2}}-ik\mu\ \ \operatorname*{\int}\limits_{0}^{\chi_{*}}d\chi\frac{e^{ik\left(1-\mu\right)\chi}}{{a(\eta-\chi)}^{2}}\right)\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (67)
@@
It is apparent from eq. (66) that, if $a(\eta_{*})\ll 1$, the second term on the right-hand side will be the dominant one of the two first terms. This is the case if the polarized radiation and the tensor mode were emitted at early times in an expanding cosmology. In the next sections, we will evaluate eq. (66) for two different backgrounds: A flat Minkowski background and an expanding, conformally flat background.
@@ VIII.2 Gravitational Polarization Sensitivity $\sigma^{s}$
Next, let us evaluate the gravitational polarization sensitivity $\sigma^{s}(\hat{k},\hat{p})$ of eq. (64). It is defined as a purely geometric quantity in terms of two spatial directions $\hat{k}$ and $\hat{p}$ and the tensor polarization basis $\sigma_{\mu\nu}^{s}$. First, the tensor basis $\sigma_{\mu\nu}^{s}(\hat{k})$ for a given direction of propagation ${\hat{k}}^{\mu}$ must be chosen. With the choice of an orthonormal polarization basis $e_{A}^{\mu}(\hat{k}),A=1,2$ perpendicular to the propagation direction $\hat{k}$, the basis tensor $\sigma_{\mu\nu}^{s}$ can be defined by
@@
[…]
$\sigma^{+}(\hat{k},\hat{p})=-\sin 2\phi{\sin}^{2}\theta=-\left(1-\mu^{2}\right)\sin 2\phi$
$\sigma^{\times}(\hat{k},\hat{p})=\cos 2\phi{\sin}^{2}\theta=\left(1-\mu^{2}\right)\cos 2\phi.$
VIII.3 Wiggle Rate from a Tensor Mode - Close Emitter
Let us consider the simplest case first: An emitter of polarized electromagnetic radiation in the cosmological neighborhood, which will be referred to as a Close Emitter (CE). In this case, we can assume a flat background with $a(\eta)=\mathrm{const}=1$. When evaluating the wiggle rate of eq. (66) for this case, the result is
$\omega_{\left(\operatorname{CE}\right)}^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\frac{\left(1-e^{i\left(1-\mu\right)k\chi_{*}}\right)}{1-\mu}e^{-ik\eta}\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}.$ (72)
To lowest order, proper time of the observer equals coordinate time; $\tau=\eta+\mathcal{O}(h)$. Since $\omega_{\mathrm{CE}}^{(T)}$ is $\mathcal{O}(h)$, we may interchange proper time and coordinate time in the above expression. Then, integrating with respect to proper time $\tau$ gives the polarization angle
$\Omega_{\left(\operatorname{CE}\right)}^{\left(T\right)}(\tau)=\mathrm{Re}\left\{\left(\sigma^{+}(\hat{k},\hat{p})h_{+}^{0}(k)+\sigma^{\times}(\hat{k},\hat{p})h_{\times}^{0}(k)\right)\frac{\left(e^{ik(1-\mu)\chi_{*}}-1\right)}{\left(1-\mu\right)}e^{-ik\tau}\right\}+\overline{\Omega},$ (73)
where the integration constant $\overline{\Omega}$ is the mean angle, which in the following will be set to zero.
In order to calculate $\Omega_{(\mathrm{CE})}^{(T)}$, we need to separate the amplitude and phase of each complex factor on the right-hand side of eq. (73). First, the factor that depends on the distance $\chi_{*}$ to the emitter can be written in polar form as
@@ VIII.3 Wiggle Rate from a Tensor Mode - Close Emitter
$\psi(\mathcal{K},\hat{p})\equiv\arctan(\frac{\sigma_{+}H_{+}\sin\alpha_{+}+\sigma_{\times}H_{\times}\ \ \sin\alpha_{\times}}{\sigma_{+}H_{+}\cos\alpha_{+}+\sigma_{\times}H_{\times}\ \ \cos\alpha_{\times}}).$ (75)
@@
$\Omega_{\left(\operatorname{CE}\right)}^{\left(T\right)}(\tau)=W_{\left(\operatorname{CE}\right)}(\mathcal{K},\hat{p},\chi_{*})\sin(k\tau+\gamma_{\left(\operatorname{CE}\right)}(\mathcal{K},\hat{p},\chi_{*})),$ (76)
@@
$W_{\left(\operatorname{CE}\right)}(\mathcal{K},\hat{p},\chi_{*})\equiv 2T(\mathcal{K},\hat{p})\frac{\sin(\frac{\left(1-\mu\right)}{2}k\chi_{*})}{1-\mu}$ (77)
@@
$\gamma_{\left(\operatorname{CE}\right)}(\mathcal{K},\hat{p},\chi_{*})\equiv-\left(\frac{\left(1-\mu\right)}{2}k\chi_{*}+\psi(\mathcal{K},\hat{p})\right).$
The polarization wiggling amplitude $W_{(\mathrm{CE})}(\mathcal{K},\hat{p},\chi_{*})$ defines the polarization wiggling response of polarized electromagnetic radiation emitted at distance $\chi_{*}$ from the detector and in direction $\hat{p}$ to a plane gravitational wave with state vector $\mathcal{K}$. The amplitude $W_{(\mathrm{CE})}$ and phase $\gamma_{(\mathrm{CE})}$ vary with the state $\mathcal{K}$ of the gravitational wave as well as the position of the emitter relative to the detector. We see from eq. (76) that the polarization wiggling frequency is equal to the gravitational wave frequency.
Let us take a look at the conditions under which the amplitude $W_{(\mathrm{CE})}$ vanishes. First of all, for inclinations $\mu=\pm 1$, the direction of the emitted radiation is aligned with the direction of the gravitational tensor mode; $\hat{p}=\pm\hat{k}$. From eqs. (70) and (71) follows that $\sigma^{+}(\hat{k},\pm\hat{k})=\sigma^{\times}(\hat{k},\pm\hat{k})=0$, which implies that
$W_{\left(\operatorname{CE}\right)}(\mathcal{K},\pm\hat{k},\chi_{*})=0.$
@@
Next, there are certain emitter distances that give vanishing $W_{(\mathrm{CE})}$ for $|\mu|<1$:
$\chi_{*}=\frac{2\pi N}{\left(1-\mu\right)k},N=1,2,...$ (78)
Finally, from eq. (77) follows that
${\lim}_{k\chi_{*}\rightarrow 0}W_{\left(\operatorname{CE}\right)}(\mathcal{K},\hat{p},\chi_{*})=0.$
Thus, a detector set up to measure polarization wiggling of electromagnetic radiation emitted by an emitter at distance $\chi_{*}$ will be insensitive to gravitational radiation with wave length significantly larger than the emitter distance $\chi_{*}$.
With the exception of these cases as well as other border cases where $T(\mathcal{K},\hat{p})$ vanishes for certain combinations of gravitational wave state $\mathcal{K}$ and polarized radiation direction $\hat{p}$, $W_{(\mathrm{CE})}(\mathcal{K},\hat{p},\chi_{*})\neq 0$. Therefore, polarized electromagenetic radiation is in general subject to polarization wiggling in the presence of gravitational radiation.
VIII.4 Wiggle Rate from a Tensor Mode - Distant Emitter
Next, let us consider a distant emitter (DE), which we define as an emitter of polarized electromagnetic radiation that is cosmologically distant, with a scale factor at the time of emission that is significantly less than 1. We will assume a conformally flat background with ${a(\eta)}^{\prime}\neq 0$ and evaluate the wiggle rate of eq. (66). In order to get analytical results and gain an understanding of the results, we will make the simplifying assumption that the observed electromagnetic radiation is emitted in a matter-dominated spacetime. Although not a realistic model of our universe at later times, we expect that this approximation is a reasonable approximation for tensor modes of smaller scale as long as they are sufficiently damped at the end of the matter-dominate era to not yield significant contributions to the wiggling rate from later times. In addition, it is a useful approximation to get a qualitative understanding of how gravitational tensor modes induce polarization wiggling in an expanding universe.
It is known that a matter-dominated universe has a scale factor  $a(\eta)={(\eta/\eta_{0})}^{2}$, where $\eta_{0}$ is the time at which $a(\eta_{0})=1$. This enables us to evaluate the third term on the right-hand side of eq. (66). Reparametrizing the integral in terms of the variable $x\equiv k\eta$, we get
$ik\mu\operatorname*{\int}\limits_{\eta_{*}}^{\eta}d\eta^{\prime}\frac{e^{-ik\left(1-\mu\right)\eta^{\prime}}}{{a(\eta^{\prime})}^{2}}=ik\mu\eta^{2}J(x,x_{*}),$
@@ VIII.4 Wiggle Rate from a Tensor Mode - Distant Emitter
$J(x,x_{*})\equiv\operatorname*{\int}\limits_{x_{*}}^{x}dt\frac{e^{-i\left(1-\mu\right)t}}{t^{2}}$
and $x_{*}\equiv k\eta_{*}$. The integral $J(x,x_{*})$ can be expressed in terms of the exponential integral $E_{2}$ as follows [32]:
$J(x,x_{*})=\frac{E_{2}(i\left(1-\mu\right)x_{*})}{x_{*}}-\frac{E_{2}(i\left(1-\mu\right)x)}{x}.$
@@
${\tilde{\omega}}_{\left(\operatorname{DE}\right)}^{\left(T\right)}(\eta)=\mathrm{Re}\left\{ik\left(1+i\mu\right)\left(\frac{1}{{a(\eta)}^{2}}-\frac{{e}^{ik(1-\mu)\chi_{*}}}{{a(\eta_{*})}^{2}}\right)e^{-ik\eta}\sum\limits_{s}\sigma^{s}(\hat{k},\hat{p})h_{s}^{0}(k)\right\}$ (79)
to leading order in $1/k{\eta}_{*}$. It is evident from eq. (79) that for radiation emitted early, while $a(\eta_{*})\ll a(\eta)=\mathcal{O}(1)$, the ${a(\eta_{*})}^{-2}$ terms on the right-hand side of eq. (79) will be dominant, and the wiggle rate depends primarily on the tensor mode amplitude at emission. This confirms our earlier claim that, for a distant emitter in an expanding universe, the dominant contributions to the present wiggle rate of polarized radiation emitted at early times come from gravitational fields present at the time of emission.
We will assume that the observation happens at present times with $\eta\simeq\eta_{0}$. To lowest order, proper time $\tau$ of the observer can be expressed in terms of coordinate time as $\tau=\eta-\eta_{0}+\mathcal{O}(h)$.
Integrating to obtain the polarization angle $\Omega_{(\mathrm{DE})}^{(T)}$ in terms of proper time $\tau$ and assuming that $a(\eta_{*})\ll a(\eta)$ yields
$\Omega_{\left(\operatorname{DE}\right)}^{\left(T\right)}(\tau)=W_{\left(\operatorname{DE}\right)}(\mathcal{K},\hat{p},\chi_{*})\sin(k\tau+\gamma_{\left(\operatorname{DE}\right)}(\mathcal{K},\hat{p},\chi_{*})),$ (80)
where
$W_{\left(\operatorname{DE}\right)}(\mathcal{K},\hat{p},\chi_{*})\equiv\frac{T(\mathcal{K},\hat{p})\sqrt{1+\mu^{2}}}{{a(\eta_{*})}^{2}}$ (81)
$\gamma_{\left(\operatorname{DE}\right)}(\mathcal{K},\hat{p},\chi_{*})\equiv-\left(\xi+\left(1-\mu\right)k\chi_{*}+\psi(\mathcal{K},\hat{p})\right),$ (82)
and  $T(\mathcal{K},\hat{p})$ and $\psi(\mathcal{K},\hat{p})$ are given by eqs. (74) and (75) above, while $\xi\equiv\arg(1+i\mu)=\arccos(1/\sqrt{1+\mu^{2}})$.
We see from eq. (80) that the polarization wigging frequency is equal to the gravitational wave frequency. This is similar to the result of eq. (76) for a close emitter. The most striking difference between the results of eqs. (76) and (80) is between the amplitudes $W_{(\mathrm{CE})}$ and $W_{(\mathrm{DE})}$ and their dependency on the emitter distance $\chi_{*}$. For a close emitter, the wiggling amplitude oscillates with emitter distance, and polarization wiggling is suppressed for emitter distances that satisfy eq. (78). This is not the case for a distant emitter as the wiggling amplitude increases slowly with emitter distance. The reason for this can be found by comparing the close emitter wiggle rate $\omega_{(\mathrm{CE})}^{(T)}(\eta)$ of eq. (72) with the distant emitter wiggle rate ${\tilde{\omega}}_{(\mathrm{DE})}^{(T)}(\eta)$ of eq. (79). The factor $(1-e^{i(1-\mu)k\chi_{*}})$ of eq. (76) can be seen as giving rise to two wiggling components with the same amplitude, but with a phase difference of $(1-\mu)k\chi_{*}+\pi$. When varying the emitter distance, these two components will interfere destructively with a regular distance interval. Similarly, for a distant emitter, the corresponding factor $({a(\eta)}^{-2}-e^{i(1-\mu)k\chi_{*}}{a({\eta}_{*})}^{-2})$ of eq. (79) gives rise to two wigging components, but with very different amplitudes and therefore no interference with varying emitter distance.